Hi, I have a question. If you could help, it'd be great. Thanks
How many gallons of milk containing 1% fat is mixed with milk containing 4% fat to obtain a 15 gallon mixture containing 2% fat?
:

Answer by ankor@dixie-net.com(22740)

How many gallons of milk containing 1% fat is mixed with milk containing 4% fat to obtain a 15 gallon mixture containing 2% fat?
:
Let x = amt of 4% fat milk required to accomplish this
the total is to be 15 gal, therefore:
(15-x) = amt of 1% fat milk needed
;
A typical mixture equation using the decimal equiv
.01(15-x) + .04x = .02(15)
.15 - .01x + .04x = .30
-.01x + .04x = .30 - .15
.03x = .15
x = .15/.03
x = 5 gal of 4% milk is needed
then
15 - 5 = 10 gal of 1% milk
:
:
Check this in the original equation